0.125 m/s b, 0.25 ms Water has a density of 1000 kgm. The depth of a swimming pool at the deep end is about 50 meters. a. What is the volume of a column of water 5.0 m and m in cross-seetional area? deep 0.8 b. What is the mass of this column of water? c. What is weight of column of water in newtons? this d. What is the gauge pressure on the bottom of the poor?
Answer #1

A.

Volume = Cross-sectional area*depth

Volume = 5*0.8 = 4 m^3

B.

mass = density*Volume

mass = 1000*4 = 4000 kg

C

Weight = mass*g

Weight = 4000*9.81 = 39240 N

D.

Gauge pressure = Patm + rho*g*H

= 1.015*10^5 + 1000*9.81*5

= 1.505*10^5 Pa